OpenStudy (anonymous):

Which equation is correct? Answer sec x° = adjacent ÷ opposite cot x° = adjacent ÷ opposite cot x° = opposite ÷ adjacent sec x° = opposite ÷ adjacent

OpenStudy (anonymous):

second

OpenStudy (anonymous):

Thanks :), I worked it out, i think thats right :)

OpenStudy (anonymous):

sin=opposite/hypotenuse cos=adjacent/hypotenuse cot=cos/sin=adjacent/opposite sec=1/cos=hypotenuse/adjacent

OpenStudy (anonymous):

thank you thank you :3

OpenStudy (anonymous):

np

OpenStudy (ajprincess):

tan=opposite/adjacent cosec=1/sin=hypotenuse/opposite